OpenStudy (wcrmelissa2001):

2 questions. Really appreciate any help :D

OpenStudy (faiqraees):

Go on

OpenStudy (wcrmelissa2001):

er sorry gimme 5 minute to slowly type the question

OpenStudy (wcrmelissa2001):

\[\frac{ 1 }{ (x-1)(x-2) }+\frac{ 1 }{ (x-2)(x-3) }+...+\frac{ 1 }{ (x-n)(x-(n+1)) } where n is a posireal number\]

OpenStudy (wcrmelissa2001):

sorry it shld say when n is a positvie real number

OpenStudy (wcrmelissa2001):

2nd qn: \[\frac{ x+a }{ b }≥\frac{ a }{ x+b } where, a,b>0\]

OpenStudy (mathmale):

Good job of typing in your problem statement in Equation Editor. Please add the instructions for this problem: what are you supposed to calculate or find?

OpenStudy (wcrmelissa2001):

Sorry. QN 1: Find the sum QN2: just solve for x the normal inequalities way

OpenStudy (mathmale):

\[\frac{ 1 }{ (x-1)(x-2) }+\frac{ 1 }{ (x-2)(x-3) }+...+\frac{ 1 }{ (x-n)(x-(n+1)) }\]

OpenStudy (mathmale):

I ask you to experiment with this as follows: Separate the first and second terms into two separate fractions each, resulting in four fractions. If you're fortunate, and if you do the algebra correctly, you'll find that two of the resulting four fractions have the same denominator, and that one fraction is + and the other -. Cancel them, leaving yourself with just 2 fractions. This is called a "telescoping series." If you have a book or online materials , look up "telescoping series" for more information.

OpenStudy (wcrmelissa2001):

THANKS :D i'll try that later